# SymBoltz.jl: a symbolic-numeric, approximation-free and differentiable linear Einstein-Boltzmann solver

### v1

@@ 5 Conclusion and future potential
SymBoltz is a fresh Julia package for solving the linearized Einstein-Boltzmann equations. It relaxes all approximation switching schemes found in other codes and solves a single set of stiff equations at all times with implicit integrators, and combats the performance loss with modern and efficient implicit ODE solvers, optimized linear system solvers and (soon) sparse matrix methods. It is differentiable, so one can get accurate derivatives of any output quantity with respect to any input parameter. This enables modelers to rapidly prototype new models by straightforwardly writing down new variables and equations.
Version 0.8.1 features the metric in the conformal Newtonian gauge, General Relativity, cold dark matter, photons, baryons and RECFAST recombination, massless and massive neutrinos, the cosmological constant and $w_{0}w_{a}$ dark energy, and computes luminosity distances and matter and CMB spectra. It also has some rudimentary models for Brans-Dicke gravity, quintessence dark energy and curved geometry, but these are not complete yet. We think the modular design makes it very easy to add new models and compute other quantities.
Forward-mode automatic differentiation is well-tested, while support for reverse-mode is a future goal. This would make SymBoltz very powerful for scalar loss applications like MCMCs.
With some more work, SymBoltz could grow into a fully integrated symbolic-numeric and differentiable cosmological modeling environment. For example, numerical code generation could be extended from linear equations to non-linear $N$-body simulations in a consistent and unified framework. This could alleviate the $N$-language problem ($N\gtrsim 2$) prevalent in cosmological modeling, which glues together convenient high-level languages like Python with performant low-level languages like C and Fortran, and sometimes symbolic work in a CAS like Mathematica.
SymBoltz shows that it is possible to create a symbolic-numeric, approximation-free and differentiable Boltzmann solver with modern numerical techniques. It excels at model simplicity and ease of modification. More work is needed to make it as feature-complete and fast as CAMB and CLASS that have been refined over many years. A major goal is to compute values and derivatives of perturbation-derived spectra fast enough for use with gradient-based MCMC samplers. Implicit ODE solvers, nonlinear and linear equation solvers, sparse matrix methods and differentiability are all demanding numerical techniques. Making them work robustly and efficiently together is challenging, but the current state shows promise. SymBoltz is built on (and has contributed to) several evolving Julia packages, and will continue to grow both on its own and with improvements to its dependencies.
@@
SymBoltz is easy to install from https://github.com/hersle/SymBoltz.jl. Documentation is available there, and the code is tested with continuous integration to ensure it remains correct and up-to-date (see appendix C). Anyone is encouraged to ask questions, give feedback, open issues and contribute pull requests in the repository! We hope SymBoltz offers valuable competition and new ideas on the Boltzmann solver market.
@@ Acknowledgments
I thank Aayush Sabharwal and Christopher Rackauckas for developing and maintaining ModelingToolkit.jl and DifferentialEquations.jl and answering questions. I thank Hans A. Winther for helpful suggestions, testing the code and giving feedback on it and this manuscript. I thank Julien Lesgourgues, Thomas Tram and others for developing CLASS, which has inspired SymBoltz.
@@ Appendix A List of equations and practical implementation details
This appendix summarizes the equations that define the standard $\mathrm{\Lambda CDM}$ model in SymBoltz and comments on their practical implementation. We hope this can be a useful reference for others. The list is structured like SymBoltz with one component per subsection. Variables are in units where $G=c=H_{0}=1$? and temperatures are in $\mathrm{K}$, unless otherwise is stated. These units are chosen because $G$, $c$ and $H_{0}$ can be divided out from the Einstein equations as natural units, but this requires some conversion in the recombination equations that depend explicitly on $H_{0}$. In other words, times are in units of $1/H_{0}$, distances in $c/H_{0}$ and masses in $c^{3}/H_{0}G$. When $G$, $c$ or and $H_{0}$ appear explicitly in the equations, they are in SI units and used only to convert from SI units into the dimensionless units. The equations closely follow the conventions in the seminal paper by Ma & Bertschinger (1995) and very closely matches the source code of SymBoltz with Unicode characters. It is far outside the scope of this paper to derive the equations and explain the meaning of every variable.
The independent variable is conformal time $\tau$, and all derivatives $^{\prime}=\mathrm{d}/\mathrm{d}\tau$ are with respect to it (in units of $1/H_{0}$). By default, integration starts from the early time $\tau=\tau_{i}=10^{-6}$, and is terminated when the scale factor crosses $a=1$, and the corresponding time $\tau=\tau_{0}$ is labeled today.
@@ A.4 Baryons $(b)$
Baryons are also non-relativistic, but interact with photons through Compton scattering and are subject to recombination physics. This significantly complicates their behavior. SymBoltz currently implements equations from RECFAST [^12] https://www.astro.ubc.ca/people/scott/recfast.html version 1.5.2 (Seager et al., 1999, 2000; Wong et al., 2008; Scott & Moss, 2009):
@@
Unlike other RECFAST implementations, SymBoltz does not approximate the stiff Peebles equations at early times by Saha approximations (although $X_{\mathrm{He}}^{++}$ is given by a Saha equation at all times). This is not necessary with a good implicit ODE solver. SymBoltz sets $C_{\mathrm{H}}=C_{\mathrm{He}_{1}}=1$ when $X_{\mathrm{e}}\gtrsim 0.99$ to avoid numerical instability at early times. Atomic calculations are done in SI units and converted to SymBoltz’ dimensionless units by factors of $H_{0}$ in SI units. The differential equation for $T_{b}^{\prime}$ is very stiff and sensitive to $T_{b}-T_{\gamma}$, but $T_{b}\approx T_{\gamma}$ in the early universe, so we rewrite it to a more stable differential equation for $\Delta T^{\prime}=T_{b}^{\prime}-T_{\gamma}^{\prime}$ instead, initialize $\Delta T=0$ and observe $T_{b}=\Delta T+T_{\gamma}$. The optical depth $\kappa(\tau)=\int_{\tau_{0}}^{\tau}\kappa^{\prime}(\tau^{\prime})\mathrm{d}\tau^{\prime}$ is really a line-of-sine integral into the past, but is integrated together with the background ODEs by initializing $\kappa(\tau_{i})=0$ to an arbitrary value, integrating the differential equation for $\kappa^{\prime}$ and subtracting the final value of $\kappa(\tau_{0})$ (i.e. $\int_{\tau_{0}}^{\tau}=\int_{\tau_{0}}^{\tau_{i}}+\int_{\tau_{i}}^{\tau}=\int_{\tau_{i}}^{\tau}-\int_{\tau_{i}}^{\tau_{0}}$). There is no tight-coupling approximation.
@@ A.7 Massive neutrinos $(h)$
On the other hand, massive neutrinos have intermediate masses that fall between the non-relativistic and ultra-relativistic limits. Integrals over their distribution function must be computed numerically. This is very expensive if done naively, and it is extremely important to choose a quadrature scheme that minimizes the number of sampled points. Fortunately, the momentum integrals have a structure that can be exploited: they are all in the form weighted form $I[g(x)]=\int_{0}^{\infty}\mathrm{d}x\,x^{2}f(x)g(x)$, where $f(x)=1/(e^{x}+1)$ is the equilibrium distribution function and $g(x)$ is an arbitrary function of the dimensionless momentum $x=pc/k_{B}T$ (see Ma & Bertschinger (1995) for more details). One can generally approximate $I[g(x)]\approx\sum_{i}W_{i}\,g(x_{i})$ with a weighted quadrature scheme with points $x_{i}$ and weights $W_{i}$ (more on this after the equations). In other words, the integral operator $\int_{0}^{\infty}\mathrm{d}x\,x^{2}f(x)$ is effectively replaced by the discrete summation operator $\sum_{i}\!W_{i}$ for some weights $W_{i}$.
@@
CAMB (Lewis, 2025) and CLASS (Lesgourgues & Tram, 2011) apply similar weighted quadrature strategies. They also get away with only a handful of sampled momenta, but the precise details of the computation differ slightly. For reference, here are points and weights computed by SymBoltz for $1\leq N\leq 8$ momenta:
@@ A.11 CMB power spectrum and line-of-sight integration
As first suggested by Seljak & Zaldarriaga (1996), this approach enables cheap computation for any $l$ after integrating the perturbation ODEs with only a few $l\leq l_{\text{max}}$. This drastically speeds up the computation over including all $l$ in an enormous set of coupled perturbation ODEs. SymBoltz performs the integrals with the trapezoid method using the substitution $u(\tau)=\tanh(\tau)$, which adds more points in the early universe when sampled uniformly, using $768$ points by default. Here $j_{l}$ are the spherical Bessel functions of the first kind. SymBoltz is not yet generalized to non-flat geometries, where they are replaced by hyperspherical functions. The cross-correlated angular spectrum between $\mathrm{A},\mathrm{B}\in\{\mathrm{T},\mathrm{E}\}$ is then computed from
@@ Appendix B Precision parameters for CLASS
When comparing results to CLASS in section 3, CLASS is configured with the following non-default precision parameters: {codebox}
background_Nloga = 6000tight_coupling_trigger_tau_c_over_tau_h = 1e-2tight_coupling_trigger_tau_c_over_tau_k = 1e-3radiation_streaming_approximation = 3ur_fluid_approximation = 3ncdm_fluid_approximation = 3We also set l_max_g, l_max_pol_g, l_max_ur, l_max_ncdm to the same $l_{\text{max}}$ used by SymBoltz’ model. These settings disable as many approximations as possible and reduces the impact of the tight-coupling approximation, which cannot be disabled. Oddly, we find that the parameter background_Nloga must be decreased from the default value 40000 to make the derivatives in fig. 4 stable. This parameter controls the number of points used for splining background functions in the perturbations. The default value of this parameter was changed from 3000 to 40000 in 2023, but we suspect that the increased density in points makes the splines susceptible to oscillations from numerical noise. These settings are important for good agreement between the Fisher forecasts in section 3.4. We used CLASS version 3.3.1.

### v2

@@ 6 Conclusion
SymBoltz is a fresh Julia package for solving the linear Einstein-Boltzmann equations. It features a symbolic-numeric interface where models are built from symbolic equations and compiled to efficient numerical code. This lets modelers prototype new models with few lines code at a high level. It relaxes all approximation schemes found in older codes by solving a single set of stiff equations at all times with implicit integrators. This simplifies the equations and remains performant due to high-order implicit solvers with fast generated code for ODEs and analytical and sparse Jacobians. It is differentiable, so one can get accurate derivatives of any output quantity with respect to any input parameter. Output power spectra agree with CLASS to $0.1\%$ with standard precision and can be used fit to current data.
Symbolic-numericinterfaceApproximation-freenessDifferentiabilityDetect sparsity of $\boldsymbol{J}$Calculate $\boldsymbol{J}$ robustlyEasy to specify model
Figure 8: A symbolic-numeric interface, approximation-freeness and differentiability are three main features of SymBoltz that form a synergy. Differentiability computes the ODE Jacobian $\boldsymbol{J}$ accurately and efficiently; which is used by implicit solvers to integrate stiff ODEs without approximations; which makes it easier to write down equations in simple symbolic form; which is used to find the sparsity pattern of $\boldsymbol{J}$; which speeds up $\boldsymbol{J}$ and the implicit solver. This is a self-reinforcing design.
Other recent codes like PyCosmo, Bolt.jl and DISCO-EB also incorporate some of these features. SymBoltz combines them and reinforces its design with a synergy between them, as explained in fig. 8. Together, these new codes offer a simpler alternative in high-level languages that complements traditional codes like CAMB and CLASS that are historically highly tuned around approximations in low-level languages. However, new codes need more work to reach the level of features and maturity of CAMB and CLASS that have grown over many years.
SymBoltz version 0.14.1 includes the flat Newtonian gauge metric, General Relativity, cold dark matter, photons, baryons and RECFAST recombination, massless and massive neutrinos, the cosmological constant and $w_{0}w_{a}$ dark energy with scalar perturbations, distances and matter and CMB spectra. Forward-mode automatic differentiation works for single-run purposes like Fisher forecasting. Fast reverse-mode scalar loss gradients, particularly through the perturbations, is a future milestone that could challenge emulators in gradient-based MCMC sampling. No differentiable Boltzmann code is yet capable of this.
@@
SymBoltz is easy to install from https://github.com/hersle/SymBoltz.jl. Documentation [^16] https://hersle.github.io/SymBoltz.jl/ is available, and the code is tested with continuous integration to remain correct and up-to-date (see appendix C). All questions and contributions are welcome in the repository! We hope SymBoltz proves useful and that it offers valuable ideas on the Boltzmann solver market.
@@ Acknowledgments
I thank the Julia community for creating an open ecosystem of scientific packages underlying SymBoltz. I thank Aayush Sabharwal and Christopher Rackauckas for developing ModelingToolkit.jl and OrdinaryDiffEq.jl and answering questions. I thank Hans A. Winther for helpful suggestions and feedback on the code and drafts of this paper. I thank Julien Lesgourgues, Thomas Tram and others for developing and documenting CLASS.
@@ Appendix A List of equations and practical implementation details
This appendix summarizes the equations that define the standard $\mathrm{\Lambda CDM}$ model in SymBoltz and comments on their practical implementation. We hope this can be a useful reference for others. The list is structured like SymBoltz with one component per subsection. Variables are in units where $G=c=H_{0}=1$? and temperatures are in $\mathrm{K}$, unless otherwise is stated. These units are chosen because $G$, $c$ and $H_{0}$ can be divided out from the Einstein equations as natural units, but this requires some conversion in the recombination equations that depend explicitly on $H_{0}$. In other words, times are in units of $1/H_{0}$, distances in $c/H_{0}$ and masses in $c^{3}/H_{0}G$. When $G$, $c$ or and $H_{0}$ appear explicitly in the equations, they are in SI units and used only to convert from SI units into the dimensionless units. The equations closely follow the conventions in the seminal paper by Ma and Bertschinger (1995) and very closely matches the source code of SymBoltz with Unicode characters. It is far outside the scope of this paper to derive the equations and explain the meaning of every variable.
The independent variable is conformal time $\tau$, and all derivatives $^{\prime}=\mathrm{d}/\mathrm{d}\tau$ are with respect to it (in units of $1/H_{0}$). This is perhaps the most common parametrization in the literature, and is natural because most equations are autonomous with respect to $\tau$ (i.e. $f(\tau,\boldsymbol{u})\rightarrow f(\boldsymbol{u})$, except some multipole truncation schemes that use $1/(k\tau)$, but these are unphysical). By default, integration starts from the early time $\tau=\tau_{i}=10^{-6}$, and is terminated when the scale factor crosses $a=1$, and the corresponding time $\tau=\tau_{0}$ is labeled today.
@@ A.4 Baryons $(b)$
Baryons are also non-relativistic, but interact with photons through Compton scattering and are subject to recombination physics. This significantly complicates their behavior. SymBoltz currently implements equations from RECFAST [^17] https://www.astro.ubc.ca/people/scott/recfast.html version 1.5.2 (Seager et al., 1999; Wong et al., 2008; Scott and Moss, 2009):
@@
Unlike other RECFAST implementations, SymBoltz does not approximate the stiff Peebles equations at early times by Saha approximations (although $X_{\mathrm{He}}^{++}$ is given by a Saha equation at all times). This is not necessary with a good implicit ODE solver. SymBoltz sets $C_{\mathrm{H}}=C_{\mathrm{He}_{1}}=1$ when $X_{\mathrm{e}}\gtrsim 0.99$ to avoid numerical instability at early times. Atomic calculations are done in SI units and converted to SymBoltz’ dimensionless units by factors of $H_{0}$ in SI units. The differential equation for $T_{b}^{\prime}$ is very stiff and sensitive to $T_{b}-T_{\gamma}$, but $T_{b}\approx T_{\gamma}$ in the early universe, so we rewrite it to a more stable differential equation for $\Delta T^{\prime}=T_{b}^{\prime}-T_{\gamma}^{\prime}$ instead, initialize $\Delta T=0$ and observe $T_{b}=\Delta T+T_{\gamma}$. The optical depth $\kappa(\tau)=\int_{\tau_{0}}^{\tau}\kappa^{\prime}(\tau^{\prime})\mathrm{d}\tau^{\prime}$ is really a line-of-sight integral into the past, but is integrated together with the background ODEs by initializing $\kappa(\tau_{i})=0$ to an arbitrary value, integrating the differential equation for $\kappa^{\prime}$ and subtracting the final value of $\kappa(\tau_{0})$ (i.e. $\int_{\tau_{0}}^{\tau}=\int_{\tau_{0}}^{\tau_{i}}+\int_{\tau_{i}}^{\tau}=\int_{\tau_{i}}^{\tau}-\int_{\tau_{i}}^{\tau_{0}}$). There is no tight-coupling approximation.
@@ A.7 Massive neutrinos $(h)$
On the other hand, massive neutrinos have intermediate masses that fall between the non-relativistic and ultra-relativistic limits. Integrals over their distribution function must be computed numerically. This is very expensive if done naively, and it is extremely important to choose a quadrature scheme that minimizes the number of sampled points. Fortunately, the momentum integrals have a structure that can be exploited: they are all in the form weighted form $I[g(x)]=\int_{0}^{\infty}\mathrm{d}x\,x^{2}f(x)g(x)$, where $f(x)=1/(e^{x}+1)$ is the equilibrium distribution function and $g(x)$ is an arbitrary function of the dimensionless momentum $x=pc/k_{B}T$ (see Ma and Bertschinger (1995) for more details). One can generally approximate $I[g(x)]\approx\sum_{i}W_{i}\,g(x_{i})$ with a weighted quadrature scheme with points $x_{i}$ and weights $W_{i}$ (more on this after the equations). In other words, the integral operator $\int_{0}^{\infty}\mathrm{d}x\,x^{2}f(x)$ is effectively replaced by the discrete summation operator $\sum_{i}\!W_{i}$ for some weights $W_{i}$.
@@
CAMB (Lewis, 2025) and CLASS (Lesgourgues and Tram, 2011) apply similar weighted quadrature strategies. They also get away with only a handful of sampled momenta, but the precise details of the computation differ slightly. For reference, here are points and weights computed by SymBoltz for $1\leq N\leq 8$ momenta:
@@ A.11 CMB power spectrum and line-of-sight integration
As first suggested by Seljak and Zaldarriaga (1996), this approach enables cheap computation for any $l$ after integrating the perturbation ODEs with only a few $l\leq l_{\text{max}}$. This drastically speeds up the computation over including all $l$ in an enormous set of coupled perturbation ODEs. SymBoltz performs the integrals with the trapezoid method using the substitution $u(\tau)=\tanh(\tau)$, which adds more points in the early universe when sampled uniformly, using $768$ points by default. Here $j_{l}$ are the spherical Bessel functions of the first kind. SymBoltz is not yet generalized to non-flat geometries, where they are replaced by hyperspherical functions. The cross-correlated angular spectrum between $\mathrm{A},\mathrm{B}\in\{\mathrm{T},\mathrm{E}\}$ is then computed from
@@
A.12 Adaptive source function refinement
SymBoltz can optionally refine source functions $S(\tau,k)$ in $k$, such as when computing the matter or CMB power spectra.
First, the source function $S_{i}(\tau)=S(\tau,k_{i})$ is found on an initial grid of wavenumbers $k_{i}$ by integrating their perturbations. The method then iterates over each interval $(k_{i},k_{i+1})$, integrates the perturbations for the center $k_{i+1/2}=(k_{i}+k_{i+1})/2$ and finds $S_{i+1/2}(\tau)=S(\tau,k_{i+1/2})$. The idea is then to compare this to a linearly interpolated source $S_{\text{int}}(\tau)=(S_{i}(\tau)+S_{i+1}(\tau))/2$ with some error measure err, for example from $\texttt{err}^{2}=\int\mathrm{d}\tau\big(S_{\text{int}}(\tau)-S_{i+1/2}(\tau)\big)^{2}$. If the interpolation is poor and err is greater than some absolute and relative tolerances, the left and right half-intervals $(k_{i},k_{i+1/2})$ and $(k_{i+1/2},k_{i+1})$ are recursively refined in the same way.
This gives $S(\tau,k)$ by integrating as few kk–modes as possible. Instead of manually creating a non-uniform $k$-grid with many hyperparameters, the process is controlled by only one tolerance. It parallelizes over each interval and can also be done in $\log k$.
@@ Appendix B Precision parameters for CLASS
When comparing results to CLASS in section 3, we use CLASS version 3.3.4 with the following precision parameters: {codebox}
tight_coupling_approximation = 5 # compromise_CLASS; cannot disable TCAtight_coupling_trigger_tau_c_over_tau_h = 0.01 # turn off TCA earliertight_coupling_trigger_tau_c_over_tau_k = 0.001 # turn off TCA earlierradiation_streaming_approximation = 3 # turn off RSAncdm_fluid_approximation = 3 # turn off NCDMFAur_fluid_approximation = 3 # turn off UFAevolver = 1 # for implicit ndf15; 0 for explicit rktol_perturbations_integration = 1.0e-6 # less noise; varied in work-precision diagramtol_ncdm_newtonian = 1.0e-5 # default sampling of massive neutrino momentabackground_Nloga = 6000 # for stable derivativesWe run CLASS with output = mPk to compute only $P(k)$, and output = tCl, pCl to compute only $C_{l}^{\text{TT}}$, $C_{l}^{\text{EE}}$ and $C_{l}^{\text{TE}}$ with the code’s most specialized and optimized paths. CLASS and SymBoltz are configured to use the same Newtonian gauge, $l_{\text{max}}$ for all species, RECFAST recombination model, $\tanh$-like reionization parametrization, $w_{0}w_{a}$ dark energy equations and cosmological parameter values. The settings above disable as many approximations as possible and reduce the impact of the tight-coupling approximation, which cannot be disabled. We had to decrease the parameter background_Nloga from the default value 40000 to make the derivatives in fig. 4 stable. This parameter controls the number of points used for splining background functions in the perturbations. Its default value was changed from 3000 to 40000 in 2023, but we suspect that the increased density in points makes the splines oscillate from numerical noise. Likewise, the tolerance for the perturbations ODEs is set one notch smaller than the default $10^{-5}$ to make the output stable. We use CLASS’ default sampling of massive neutrino momenta, as discussed in section 4.
When running CLASS with all approximations on in tables 1 and 7, the first block of parameters above are not used. In the work-precision comparison in fig. 7, we also vary tol_perturbations_integration and evolver as described in the figure.